OpenStudy (anonymous):

I always get stuck on this. I have 4x - 2y = 6 and I have to get into slope intercept form. I got this far 4x - 4x - 2y = 6 - 4x = 2y = 6 - 4x, now what?

OpenStudy (rosedewittbukater):

Slope intercept form is y=mx+b. m is the slope, and b is the y intercept. 4x - 2y = 6 -2y = -4x + 6 Subtract 4x from both sides y = 2x - 3 Divide both sides by -2 to get y by itself That's it. y = 2x - 3

OpenStudy (anonymous):

Thanks, you. The dividing part always confuses me because of the x.

OpenStudy (mathstudent55):

Solve the equation for y 4x - 2y = 6 Since you need to get rid of 4x from left, subtract 4x from both sides. (Write the right side as written below, -4x + 6, not 6 - 4x, to follow the y = mx + b form). 4x - 4x - 2y = -4x + 6 The 4x and -4x terms on left side add up to zero -2y = - 4x + 6 Now divide both sides by -2 y = 2x - 3

OpenStudy (anonymous):

Thanks

OpenStudy (mathstudent55):

wlcm